### v1

# Suppression of Fast Flavor Conversion by Red Turbulence in Supernovae

Abstract
Fast flavor conversions (FFCs) in supernovae, driven by neutrino-neutrino refraction, can catastrophically equilibrate flavors and potentially affect the neutrino-driven explosion. We present a pivotal insight: matter density fluctuations characterized by red spectra ($\nu<0$), naturally arising in stratified supernova environments, can suppress such instabilities by inducing accelerated decoherence. By deriving exact analytical solutions for two-flavor evolution in red turbulent matter—where correlations grow as $t^{|\nu|}$—we uncover a novel acceleration of coherence loss. This dynamical decoherence mechanism raises an effective energy barrier against the collective growth of flavor instabilities. Translating our master-equation results into an effective damping rate for FFC linear analysis, we find that realistic red turbulence ($\nu\sim-1$, fluctuation strength $\xi_{\nu}\sim 0.1$) can elevate the FFC threshold by a factor of $\sim 3-5$, potentially stabilizing regions that would otherwise undergo explosion-killing flavor equilibration (or vice versa). Our work provides the first analytical criterion for FFC suppression in turbulent media and identifies red turbulence as a critical, physics-grounded ingredient missing from current supernova models.
Introduction.—Core-collapse supernova explosions are powered by a delicate balance: neutrino heating in the gain region must narrowly exceed cooling to revive the stalled shock. Modern simulations reveal this balance is often marginal, with net heating surpassing cooling by mere percentages [4, 10]. This fine-tuning renders the explosion mechanism exquisitely sensitive to microphysical processes that alter neutrino fluxes, particularly flavor transformations.
While matter-induced (MSW) oscillations occur predominantly outside the gain region, the collective oscillations driven by neutrino-neutrino interactions present a more immediate and disruptive threat. Fast flavor conversions (FFCs), which can develop on centimeter scales within nanoseconds, rapidly equilibrate $\nu_{e}$ and $\nu_{x}$ populations. This equilibration can reduce the neutrino heating efficiency by 20–30%, thereby quenching explosions that would otherwise be marginally successful [3]. The impact of FFCs on supernova dynamics, however, remains an active area of investigation, with recent studies suggesting outcomes that range from strongly suppressive to modestly supportive depending on the progenitor and modeling details [8, 14, 15]. Linear stability analyses suggest FFCs may be ubiquitous in supernova cores [7], raising a critical question: what physical mechanisms can suppress these instabilities?
Turbulence, generated by convection and hydrodynamical instabilities, permeates supernovae and affects neutrino propagation. Beginning with Loreti and Balantekin’s master-equation approach [12], studies have examined white-noise and Kolmogorov-type spectra. Intriguingly, some turbulent spectra may even enhance FFC growth [2], deepening the theoretical challenge.
However, supernova turbulence in stratified regions—like the proto-neutron star atmosphere or convective shells—differs fundamentally. It exhibits ”red spectra” with negative spectral exponents $\nu<0$, where temporal correlations grow algebraically ($K(t)\sim t^{|\nu|}$) at short times. This red turbulence reflects persistent, coherent structures rather than a disordered cascade and is a natural outcome in density-stratified environments [5]. Its impact on neutrino flavor evolution has remained unexplored.
In this Letter, we present exact analytical solutions for neutrino evolution in red turbulent matter. Solving the non-Markovian master equation with kernel $K(t)=\xi_{\nu}(t/\tau_{c})^{|\nu|}e^{-t/\tau_{c}}$ ($\nu<0$), we discover that red turbulence generates fractional integral operators of order $1+|\nu|$, contrasting with the fractional derivatives of Kolmogorov turbulence. This leads to a key physical effect: accelerated decoherence—slower damping of flavor coherence compared to white noise of equivalent fluctuation strength.
We translate this microphysical insight to FFCs: the derived decoherence rate $\Gamma_{\text{dec}}$ acts as an effective damping term in the linear stability analysis. For realistic supernova parameters ($\nu\sim-1$, $\xi_{\nu}\sim 0.1$, $\tau_{c}\sim 0.1$ km), $\Gamma_{\text{dec}}\sim\mathcal{O}(10^{3})\,\text{km}^{-1}$ can raise the FFC threshold by a factor of 3–5, potentially stabilizing regions vulnerable to explosion-killing flavor equilibration or vice versa. Our work provides both a novel analytical framework for neutrinos in correlated media and a physics-based mechanism to resolve the tension between predicted FFC instabilities and the requirement of sustained neutrino heating for successful explosions.
Analytical Two-Flavor Solutions for Red Turbulence.—We present exact analytical solutions for neutrino evolution in turbulent matter with red spectra ($\nu<0$), where density correlations grow algebraically as $t^{|\nu|}$. The complete mathematical derivation is provided in the companion paper [1].
All quantities are expressed in natural units ($c=\hbar=1$). For red turbulence with spectral exponent $\nu<0$, the correlation function exhibits the distinctive form $K(t)\propto t^{-\nu}e^{-t/\tau_{c}}=t^{|\nu|}e^{-t/\tau_{c}}$, contrasting with the decaying correlations ($t^{-5/3}$) of Kolmogorov turbulence. The fundamental parameters are:
Table 1: Fundamental parameters for neutrino propagation in red turbulent matter ($\nu<0$, natural units $c=\hbar=1$).
Symbol Description Dimension
$\nu$ Spectral exponent ($\nu<0$ for red turbulence) dimensionless
$L_{c}$ Turbulence correlation scale $E^{-1}]$
$\tau_{c}=L_{c}$ Correlation time $E^{-1}]$
$|\nu|=-\nu$ Growth exponent of correlations dimensionless
$G_{F}$ Fermi constant $E^{-2}]$
$E$ Neutrino energy $[E]$
$\Delta m^{2}$ Neutrino mass-squared difference $E^{2}]$
$\theta$ Vacuum mixing angle dimensionless
$\xi_{\nu}$ Red turbulence strength parameter dimensionless
$\epsilon$ Dissipation scale cutoff $E^{-1}]$
The non-Markovian master equation for neutrino evolution in red turbulent matter is:
$\frac{d}{dt}\langle\hat{\rho}(t)\rangle=$
$-i[\hat{H}_{0},\langle\hat{\rho}(t)\rangle]-g_{\nu}\int_{0}^{t}ds\,(t-s)^{|\nu|}e^{-(t-s)/\tau_{c}}[\sigma_{z},[\sigma_{z},\langle\hat{\rho}(s)\rangle]],$ (1)
where $\hat{H}_{0}=\frac{\Delta_{m}}{2}(-\cos 2\theta_{m}\sigma_{z}+\sin 2\theta_{m}\sigma_{x})$ is the effective Hamiltonian in the matter basis, with $\Delta_{m}=\sqrt{(\frac{\Delta m^{2}}{2E}\cos 2\theta-\sqrt{2}G_{F}n_{e})^{2}+(\frac{\Delta m^{2}}{2E}\sin 2\theta)^{2}}$ being the matter oscillation frequency. The kernel $K(t)=g_{\nu}t^{|\nu|}e^{-t/\tau_{c}}$ encodes the algebraic growth characteristic of red turbulence, with $g_{\nu}$ quantifying the turbulence strength [1].
The exact Laplace-space solution for the electron neutrino survival probability is:
$\tilde{P}_{ee}(z)=\frac{z[z+4\tilde{K}_{\nu}(z)]+\Delta_{m}^{2}\cos^{2}2\theta_{m}}{z\left\{[z+4\tilde{K}_{\nu}(z)]^{2}+\Delta_{m}^{2}\right\}},$ (2)
where $\tilde{K}_{\nu}(z)=\int_{0}^{\infty}e^{-zt}K(t)dt$ is the Laplace transform of the red turbulence kernel.
For red turbulence with $\nu<0$:
$\tilde{K}_{\nu}(z)=g_{\nu}\Gamma(1-\nu)(z+\tau_{c}^{-1})^{\nu-1},$ (3)
where $\Gamma(1-\nu)>0$ for $\nu<0$. The key mathematical distinction emerges: while Kolmogorov turbulence ($\nu=5/3$) generates fractional derivatives of order $2/3$, red turbulence ($\nu<0$) produces fractional integrals of order $1+|\nu|$. This fundamentally alters the decoherence dynamics.
The perturbative approach naturally extends to three flavors via Gell-Mann matrix generalizations [9].
Asymptotic Analysis for Red Turbulence.—Red turbulence produces accelerated decoherence with distinctive scaling laws. We define dimensionless parameters: turbulence strength $\xi_{\nu}\approx(\delta n_{e}/\bar{n}_{e})^{2}$, memory parameter $\alpha=\Delta_{m}\tau_{c}$, and scaled time $x=\Delta_{m}t$. The survival probability exhibits universal scaling:
$P_{ee}(x)=\frac{1}{2}+\frac{1}{2}\cos 2\theta_{m}\,F_{\nu}(x,\alpha,\xi_{\nu})\cos(x),$ (4)
where $F_{\nu}$ is the red turbulence scaling function.
Regime I: Early-Time Accelerated Decoherence ($x\ll\alpha$): For $t\ll\tau_{c}$, red turbulence produces accelerated decoherence scaling as $x^{2+|\nu|}$:
$P_{ee}^{(I)}(x)=1-\frac{\sin^{2}2\theta_{m}}{4}x^{2}-\xi_{\nu}\,\mathcal{F}_{\text{red}}(x)+\mathcal{O}(x^{4+2|\nu|}),$ (5)
where:
$\mathcal{F}_{\text{red}}(x)\approx C_{\nu}x^{2+|\nu|}+C_{2}x^{3+|\nu|}+\cdots,$ (6)
with $C_{\nu}=[\Gamma(1-\nu)/(2+|\nu|)][1/(1+|\nu|)-1/(2+|\nu|)]>0$. For $\nu=-1$, this yields $x^{3}$ scaling, significantly faster than Kolmogorov’s $x^{5/3}$.
Regime II: Crossover at Correlation Time ($x\sim\alpha$): At $t\sim\tau_{c}$:
$F_{\nu}(x,\alpha,\xi_{\nu})=\exp\left[-\xi_{\nu}\Gamma(1-\nu)\frac{\gamma(1-\nu,x/\alpha)}{\Gamma(1-\nu)}\right]\left[1+\mathcal{O}(\xi_{\nu}^{2})\right],$ (7)
where $\gamma(s,z)$ is the incomplete Gamma function. Red turbulence shows enhanced memory effects due to the integral nature of the fractional operator.
Regime III: Late-Time Markovian Limit ($x\gg\alpha$): For $t\gg\tau_{c}$:
$F_{\nu}(x,\alpha,\xi_{\nu})\approx\exp\left(-\frac{\xi_{\nu}}{\alpha}x\right)\left[1+\mathcal{O}\left(\frac{e^{-x/\alpha}}{x^{1-\nu}}\right)\right],$ (8)
recovering exponential decay with rate $\Gamma_{\text{eff}}=\xi_{\nu}/\tau_{c}$, but with algebraic corrections decaying as $x^{-(1+|\nu|)}$.
Figure 1: Electron neutrino survival probability $P_{ee}(x)$ for turbulence with different spectral indices $\nu$, demonstrating the effect of temporal correlations on decoherence. White noise ($\nu=0$, blue solid line) produces the strongest decoherence, reaching the lowest asymptotic value. Red noise with $\nu=-0.5$ (orange dashed), $\nu=-1.0$ (green dash-dot), and $\nu=-2.0$ (red dotted) shows progressively weaker decoherence as $\nu$ decreases, with higher asymptotic survival probabilities. This indicates that long-range temporal correlations (red spectra) preserve flavor coherence more effectively than uncorrelated (white) fluctuations, despite having identical integrated turbulence strength $\xi=0.1$. The small-scale regularization $\epsilon=0.01$ (we introduce a physical cutoff $\epsilon\sim\tau_{\text{diss}}$ at the dissipation scale, modifying the kernel to $K(t)\propto t^{-\nu}\exp(-\epsilon/t)$. For $\nu<0$, this regularization ensures mathematical consistency while having negligible effect on physical results) ensures convergence of all memory integrals.
The key distinctions of red turbulence are: (1) Accelerated early-time scaling: $t^{2+|\nu|}$ vs. Kolmogorov’s $t^{5/3}$ (2) Fractional integrals rather than derivatives (3) Stronger memory effects at crossover (4) Faster perturbative convergence due to regular kernel at $t=0$
Mathematical Foundation.—Red turbulence generates fractional integral operators of order $1+|\nu|$, establishing a new mathematical paradigm. The regular kernel $t^{|\nu|}$ (vs. Kolmogorov’s singular $t^{-5/3}$) ensures well-behaved perturbation theory with factorial convergence $1/\Gamma(n(2+|\nu|))$. This contrasts with numerical approaches [11] that capture phase interference effects; our analytical solutions provide the dominant decoherence physics essential for understanding red turbulence signatures.
Physical Implications for Supernova Heating.—Our analytical framework provides quantitative predictions for how red turbulence affects neutrino flavor evolution in supernovae. The key quantity is the effective decoherence rate $\Gamma_{\text{dec}}$ derived from our master equation solution, which modifies the FFC linear stability criterion to:
Im[$\imaginary[\omega_{\text{FFC}}]>\Gamma_{\text{dec}}(\nu,\xi_{\nu},\tau_{c}),$ (9)
where Im[$\imaginary[\omega_{\text{FFC}}]$ is the FFC growth rate in static matter.
To quantify the potential suppression, we evaluate $\Gamma_{\text{dec}}$ using turbulence parameters from supernova simulations. Multi-dimensional studies show significant density fluctuations in the gain region, with relative amplitudes $\delta n_{e}/\bar{n}_{e}\sim 40\%$ yielding $\xi_{\nu}\sim 0.16$ [13]. The turbulence correlation scale in stratified regions like the proto-neutron star atmosphere is typically $L_{c}\sim 0.1-1$ km, giving $\tau_{c}\sim 0.1$ km in natural units [6]. For stratified turbulence, spectral exponents $\nu$ range from 1.0-1.0 (strong stratification) to 0.5-0.5 (moderate stratification) [5].
For the representative case $\nu=-1$, our solution gives $\Gamma_{\text{dec}}\sim\xi_{\nu}\tau_{c}\omega^{2}$ when $\omega\tau_{c}\ll 1$, where $\omega$ characterizes the neutrino oscillation frequency. In the gain region, the neutrino-neutrino potential $\mu_{\nu\nu}\sim 10^{2}-10^{4}$ km-1. Taking $\omega\sim\mu_{\nu\nu}\sim 10^{3}$ km-1 as a typical value, we obtain:
$\Gamma_{\text{dec}}\sim 0.16\times 0.1\,\text{km}\times(10^{3}\,\text{km}^{-1})^{2}\sim 1.6\times 10^{4}\,\text{km}^{-1}.$ (10)
This damping rate exceeds typical FFC growth rates in the gain region ($\sim 10^{1}-10^{3}$ km-1) by approximately an order of magnitude. The instability threshold is thus raised by a factor of 3-10, with the conservative estimate being a 3-5 fold enhancement.
The implications for explosion dynamics stem from the delicate heating-cooling balance. Simulations show that successful explosions typically require net positive heating with $Q^{+}>Q^{-}$ by only a few percent [4]. Turbulence itself can aid explosion by enhancing neutrino heating [6], but FFCs threaten to counteract this by reducing the effective $\nu_{e}$ and $\bar{\nu}_{e}$ fluxes. Our mechanism suggests that the same stratified turbulence that enhances heating may also suppress FFCs through decoherence.
While quantifying the exact impact on explosion outcomes requires incorporating our suppression criterion into full simulations, the magnitude of the threshold enhancement suggests it could be significant. In marginal cases where small changes in heating determine success or failure, suppressing FFCs could tip the balance toward explosion. This is particularly relevant given that FFCs are predicted to be most prevalent in regions where red turbulence naturally occurs [5].
For practical implementation, we propose that future supernova simulations extract turbulence parameters ($\nu$, $\xi_{\nu}$, $\tau_{c}$) along neutrino trajectories and apply the criterion Im[$\imaginary[\omega_{\text{FFC}}]>\Gamma_{\text{dec}}$ to determine FFC viability region-by-region. This provides a physics-based approach to assess whether and where FFCs might develop in realistic turbulent environments.
Conclusions.—We have developed the first analytical framework for neutrino flavor evolution in matter with red turbulent spectra, characterized by negative spectral exponents $\nu<0$. Our exact solutions to the non-Markovian master equation reveal that red turbulence generates fractional integral operators of order $1+|\nu|$, fundamentally distinguishing it from conventional turbulence models that produce fractional derivatives.
The key physical discovery is that red turbulence induces accelerated decoherence: for the same integrated fluctuation strength $\xi_{\nu}$, flavor coherence persists longer than in white-noise environments but decays differently than in Kolmogorov-type turbulence. This decoherence behavior emerges from the algebraically growing temporal correlations $K(t)\sim t^{|\nu|}$ characteristic of stratified astrophysical plasmas.
Applying our results to the supernova fast flavor conversion problem, we find that the derived decoherence rate $\Gamma_{\text{dec}}(\nu,\xi_{\nu},\tau_{c})$ can act as an effective damping mechanism in the linear stability analysis. For realistic parameters in supernova gain regions and proto-neutron star atmospheres ($\nu\sim-1$, $\xi_{\nu}\sim 0.1$, $\tau_{c}\sim 0.1$ km), this damping rate is $\mathcal{O}(10^{3}-10^{4})$ km-1—comparable to or exceeding typical FFC growth rates. Consequently, red turbulence can raise the FFC instability threshold by a factor of 3-5, potentially suppressing explosion-killing flavor equilibration in critical regions.
Our work establishes several important advances: 1. Mathematical foundation: Exact solutions for neutrino propagation in red turbulent media, expressed through Mittag-Leffler functions and fractional calculus. 2. Physical mechanism: Identification of accelerated decoherence as a novel effect of long-range temporal correlations. 3. Astrophysical application: A quantitative criterion for FFC suppression that can be implemented in supernova simulations. 4. Paradigm shift: Red turbulence as a self-consistent suppressor of flavor instabilities, rather than merely a source of noise.
These results suggest that the stratified turbulence naturally present in supernovae may protect the neutrino heating mechanism from disruptive fast flavor conversions. This could provide a resolution to the tension between the linear stability prediction of ubiquitous FFCs and the requirement of sustained $\nu_{e}$ dominance for successful explosions, although it is also possible that it exacerbates [8, 14, 15] the problem.
Looking forward, our analytical framework provides essential benchmarks for numerical simulations and opens new directions for studying quantum systems in correlated astrophysical environments. The connection between negative spectral exponents, fractional integrals, and decoherence dynamics represents a fundamental advance with applications beyond supernovae to neutron star mergers, accretion disks, and early universe cosmology.
Companion paper
Detailed derivations, parameter mapping to physical turbulence, and numerical validation are provided in the companion paper [1].
Acknowledgements.
We thank Thierry Foglizzo for insightful discussions and the developers of the mpmath Python library for high-precision numerical evaluation of special functions. During the preparation of this work, the authors used DeepSeekV3 to improve readability and language. After using this tool, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication. This work is supported by the Fundamental Research Funds for the Central Universities under No. 020114380057, and by K. C. Wong Educational Foundation. A.A. is supported by the National Science Foundation of China (NSFC) through the grant No. 12350410358; the Talent Scientific Research Program of College of Physics, Sichuan University, Grant No. 1082204112427; the Fostering Program in Disciplines Possessing Novel Features for Natural Science of Sichuan University, Grant No.2020SCUNL209 and 1000 Talent program of Sichuan province 2021. S.Z. is supported by the National Science Foundation of China (NSFC) through the grant Nos. 12288102, 12473031; the Yunnan Revitalization Talent Support Program–Young Talent project; and the Yunnan Fundamental Research Project No. 202501AS070078.